OpenStudy (anonymous):

Write the quadratic equation in standard form. x + 5 = x(x − 9) Write the quadratic equation in standard form 2(x − 3)2 = 1

OpenStudy (anonymous):

Quadratic Standard form is ax²+bx+c=0 For the first one, first distribute the right side. x+5=x²-9x Then move the right side to the left by subtracting x² and adding 9x, which sets it equal to zero. -x²+9x+x+5=0 Combine like terms -x²+10x+5=0 x² must be positive, so multiply all by -1 Final answer x²-10x-5=0

OpenStudy (anonymous):

How about the 2nd one?

OpenStudy (anonymous):

Second one is the second 2 supposed to be a power like x² < this 2?

OpenStudy (anonymous):

2(x-3)²=1 ?

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

Here's what I got, but I'm not 100% on this one. 2(x-3)²=1 Distribute power (2)(x²+9)=1 Then multiply by 2 2x²+18=1 Subtract 1 2x²+18-1=0 Divide by 2 x²+9-1/2=0